OpenStudy (kirbykirby):

(Sequences): Find the minimum value of n must be greater than for the following to be true: |(7+1/n)^2-49| < e for e >0(Sequences): Find the minimum value of n must be greater than for the following to be true: |(7+1/n)^2-49| < e for e >0@Mathematics

OpenStudy (kirbykirby):

\[\left| (7+1/n)^{2} - 49\right|<e, e>0\]

OpenStudy (kirbykirby):

Just as a note, I never did the epsilon/delta stuff in Cal 1 so I am quite lost with this :(

OpenStudy (agreene):

So, if I'm understanding right... you could just write this as: \[\left| (7+\frac{1}{n})^2-49 \right|<0\] since the thing has to be less than e and e must be bigger than 0... and as it happens that inequality has no solutions. Unless I'm missing something.

OpenStudy (perl):

whats the original question in complete form

OpenStudy (perl):

dont give us drips and drabs of the question, type the full complete one ;)

OpenStudy (kirbykirby):

Assuming n-> infinity, which of the following conditions represents the minimum value n must be greater than in order to satisfy: |(7+.....)|<e for any e>o

OpenStudy (kirbykirby):

a) n>\[(7+\sqrt{49+e})/e\] b) n> \[(7-sqrt{49+e})/e\] c) n> \[(7+sqrt{49-e})/e\] d) n> \[(7-sqrt{49-e})/e\]

OpenStudy (kirbykirby):

ok the sqrt is supposed to be over the 49-/+e all the time..

OpenStudy (kirbykirby):

and e here is epsilon...

OpenStudy (perl):

you want to show that lim ( 7 + 1/n) ^2 = 49

OpenStudy (perl):

as n-> oo

OpenStudy (kirbykirby):

I dunno none of my answers match... I got this: n > 7 - sqrt(e+49) ------------- 7*sqrt(e+49)